Find the absolute maximum and minimum values of f(x)=−(x2−x)23 over the interval [-4, 3].
1
Expert's answer
2020-03-05T11:36:52-0500

The given function is,

f(x) = -(x2 - x)23

and the interval is [-4,3].


Since, f(x) is a polynomial, it is continuous everywhere. Therefore, it is continuous on the given interval.


Differentiating f(x) to get the first derivative and to calculate the critical points of f(x) :


f'(x) = d(-(x2 - x)23)/dx

f'(x) = (-23(x2 - x)22) × (2x - 1)


To get the critical points, we make f'(x) equal to 0.


i.e., (-23(x2 - x)22) × (2x-1) = 0

i.e., (x2 - x)22 × (2x - 1) = 0

i.e., (x2 - x)22 = 0 or 2x - 1 = 0

x2 - x = 0 or 2x = 1

x(x - 1) = 0 or x = 1/2

x = 0 or x = 1 or x = 1/2


Therefore, we get three critical points at x = 0, x = 1/2 and x = 1, and all these points lie in the given interval.


Evaluating the function at the critical points and the end points :


f(-4) = -( (-4)2 - (-4) )23 = -(16 + 4)23 = -(20)23

f(0) = -(02 - 0)23 = 0

f(1/2) = -( (1/2)2 - (1/2) )23 = -( (1/4) - (1/2) )23 = - (-1/4)23 = - (-1)23 × (1/4)23 = (1/4)23

f(1) = -(12 - 1)23 = -(1 - 1)23 = 0

f(3) = -(32 - 3)23 = -(9 - 3)23 = -(6)23


Therefore, we see from the above list that the absolute maximum of f(x) is (1/4)23 which occurs at x = (1/2) (critical point), and the absolute minimum of f(x) is ( - (20)23 ) which occurs at x = (-4) (end point).